OpenStudy (anonymous):

Solve -3x^2+3-x=90=0 using quadratic formula

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

here are the choices:

OpenStudy (anonymous):

Since 30 is not an odd number, its fine to use the completing the square method, but i dont really like it

OpenStudy (anonymous):

\[x1,2 = \frac{ -30\pm \sqrt{30^2-4\times(-3)\times(-90)} }{ 2\times(-3) }\]

OpenStudy (anonymous):

i got it, but im stuck on the last step

OpenStudy (anonymous):

\[x1,2 = 5 \pm \frac{ \sqrt{-180} }{ -6 } \]

OpenStudy (anonymous):

wait why the 5?

OpenStudy (anonymous):

its supposed to be a 30

OpenStudy (anonymous):

oh i get it

OpenStudy (anonymous):

now we can insert the -6 into the square root - it becomes 36: \[x1,2 = 5 \pm \sqrt{\frac{ -180}{ 36 }}\]

OpenStudy (anonymous):

now -180/36 = -5 so square root of -5 is i times square root of 5

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

ok its B thanks

OpenStudy (anonymous):